OpenStudy (anonymous):

Find the indicated quantities for f(x)=3x^2 The average rate of change of f(x) if x changes from 1 to 4

OpenStudy (anonymous):

\[\frac{f(4)-f(1)}{4-1}\] is what you need to compute

OpenStudy (anonymous):

\[f(4)=3\times 4^2=48, f(1)=3\times 1^2=3\] you get \[\frac{48-3}{3}=\frac{45}{3}=15\]

OpenStudy (unklerhaukus):

\[\Delta f=f(x_1)-f(x_0)\]\[\Delta x=x_1-x_0\] The average rate of change is \(\frac{\Delta f(x)}{\Delta x}\)